Hi please help me with this
X-1/3 + x+2/6=6
Note that / is upon

Answer by checkley75(3666)

X-1/3+X+2/6=6
2X=6+1/3-2/6
2X=(36+2-2)/6
2X=36/6
2X=6
X=6/2
X=3
PROOF
3-1/3+3+2/6=6
3-1/3+3+2/6=6
6-1/3=6
6=6